a car travels at 80 kilometers per hours for 40 kilometers, 72 kilometers per hours for 48 kilometers and then 96 kilometers per hours for 32 kilometers. what is the average rate of speed over the entire trip
Answer by CharlesG2(834)

a car travels at 80 kilometers per hours for 40 kilometers, 72 kilometers per hours for 48 kilometers and then 96 kilometers per hours for 32 kilometers. what is the average rate of speed over the entire trip
D distance = R rate * T time
40 * km = 80 * km/hr * T1
40/80 * km * hr/km = T1
1/2 hr = T1
48 * km = 72 * km/hr * T2
48/72 * km * hr/km = T2
2/3 hr = T2
32 * km = 96 * km/hr * T3
32/96 * km * hr/km = T3
1/3 hr = T3
T = T1 + T2 + T3 = 1/2 + 2/3 + 1/3 = 1 1/2 = 3/2 = 1.5 hrs
D = D1 + D2 + D3 = 40 + 48 + 32 = 120 km
R = D/T = 120/1.5 = 80 km/hr average rate of speed